# A Survey of Text Watermarking in the Era of Large Language Models

### v3

@@ 5.4.2. Unforgeability in Publicly Detectable Scenario
Evaluating the unforgeability of watermark algorithms under publicly detectable scenarios is far more challenging, as the detection methods are open, allowing attackers to more easily attempt to crack or remove the watermark. In such scenarios, attackers could still employ previously mentioned statistical attack methods, such as spoofing attacks (Sadasivan et al., 2023), to analyze and extract the watermark’s generation rules. Since the detector is public in this case, there might be more data available for such attacks. Additionally, as watermark generation and detection algorithms are often closely related, attackers might directly analyze how the watermark detector is implemented and how the watermark is generated.
Currently, most watermarking algorithms utilize the watermark generator to discern watermark features in text details, thereby exposing the watermark generator. For example, Kirchenbauer et al. (2023a) still requires determining whether each token is on the green list during the watermark detection process, which exposes the watermark generation method. Consequently, these watermarking algorithms lack unforgeability in publicly detectable scenarios. A text watermarking algorithm with unforgeability must ensure that the watermark detector does not reveal information about the watermark generator. For instance, some current methods employ neural networks for watermark detection (Liu et al., 2023b; Abdelnabi and Fritz, 2021), achieving effective detection while preventing further information disclosure due to the black-box nature of neural networks.
@@ 6. Application for Text Watermarking
In the previous sections, we have comprehensively detailed the implementation methods of text watermarking technologies in the era of large models and how to thoroughly test these watermarking technologies. This section continues to discuss the application of text watermarking technologies in real-world scenarios. Specifically, we will primarily focus on the application of text watermarking technologies in three key areas: copyright protection (Zhao et al., 2023b, 2022; He et al., 2022a, b; Peng et al., 2023; Li et al., 2023; Tang et al., [n. d.]; Chu et al., 2023; Tang et al., 2023; Sun et al., 2023; Wang et al., 2023a; Taleby Ahvanooey et al., 2018; Mir, 2014; Iqbal et al., 2019), fake news detection (Megías et al., 2021, 2022), and academic integrity (Vasilatos et al., 2023; Zhao et al., 2022; Perkins, 2023). Firstly, we will discuss how to protect the copyright of large models and text/data sets using text watermarking technology to prevent infringement and misuse of intellectual property. Secondly, the article will explore the role of this technology in identifying and combating the spread of false information, particularly in the current domains of social media and news dissemination. Lastly, we delve into the significance of text watermarking technology in maintaining academic integrity, including its use in plagiarism detection and ensuring the originality of academic works. Meanwhile, we also provide a more illustrative depiction of the different applications of text watermarking in Figure 7.
@@ 6.1.1. Text/Dataset Copyright
Text copyright refers to the legal protection of original textual content, such as writings and online posts, ensuring unique rights for the creators. Copyrighted texts should provide sufficient information to identify their creators and sources. This protection extends beyond traditional publications like books and journals to digital-era web articles, blog posts, and other online contents. Current explorations in textual copyright primarily focus on format-based watermarking algorithms (section 3.1). The main reason is that only format-based watermarking does not require modifications to the text content, whereas other watermarking techniques necessitate changes, which may be unacceptable to some content creators.
For instance, Taleby Ahvanooey et al. (2018) utilizes layout features (e.g., spacing between words and lines) and formatting (e.g., text color, font, and height) to embed watermarks in the text. To enhance the protection of text copyrights in web content, Mir (2014) proposed an invisible digital watermark based on the textual information contained in web pages. This watermark utilizes predefined semantic and syntactic rules, which are encrypted and transformed into spaces using binary control characters, then embedded within the webpage. The watermark is embedded using the structure of HTML (Hypertext Markup Language) as a cover file. Moreover, for certain specific text formats, distinct watermark designs may be applied. For instance, Iqbal et al. (2019) have focused on embedding watermarks in MS-Word documents by utilizing unique attributes like variables, bookmarks, and ranges for copyright protection.
Although current text copyright protection methods predominantly use format-based watermarking, the advancement of large language models suggests that applying watermark algorithms with large language models to text copyright protection is an important direction for future research.
With the rise and widespread application of deep learning technology, the dataset copyright has become particularly important, which means protecting datasets from unauthorized use has emerged as a crucial issue. In the realm of dataset copyright protection, text watermarking technology plays a pivotal role. By adding watermarks to some data within a dataset, a watermarked dataset is created. Models trained on this watermarked dataset will possess detectable characteristics that prevent unauthorized use.
@@ 6.1.2. LLM Copyright
In the domain of copyright protection for large language models (LLMs), the primary objective is to safeguard these models from threats known as extraction attacks. These attacks involve extracting substantial amounts of data from large language models to train a new model. To protect LLM copyrights, a common method is to embed watermarks in the LLM’s output. This results in attackers using datasets with watermarks for training, leading to the new model’s outputs also bearing watermark characteristics. This process bears some resemblance to dataset copyright, except that in this case, the watermarked dataset is generated by a LLM. Current efforts have developed watermark algorithms for various LLM types, including embedding (Peng et al., 2023), generative (He et al., 2022a, b; Zhao et al., 2023b), and classification (Zhao et al., 2022) LLMs. The input of an embedding LLM is text, and its output is the corresponding embedding of that text. The generative LLM is currently the most commonly used LLM, with both its input and output being text. In the case of a classification LLM, the input is text, and the output is a specific category.
@@ 6.3. Fake News Detection
The application of large language models has raised two main concerns: the generation of false information due to their text-generating capabilities, and the rapid dissemination of such false information. Firstly, the powerful text-generating capability of LLMs makes them an effective tool for producing fake information. These models can rapidly generate texts that appear authentic and accurate, but may actually contain erroneous or misleading information. Owing to the high credibility and realism of the generated texts, they can be easily utilized to construct fake news or false narratives, misleading the public and distorting facts. The second concern is the rapid spread of false information. Fake information generated by LLMs, due to its high degree of realism, can spread swiftly on social media and other digital platforms, leading to the proliferation and reinforcement of incorrect viewpoints (DiResta, 2020; Zhou et al., 2023). This spread not only amplifies the impact of false information but can also lead to public confusion and distrust towards authoritative information sources. Therefore, the automatic identification of news generated by LLMs is essential.
The current research has explored utilizing watermark technology to detect fake images and videos generated by AI models, aiding in the identification of fake news. For instance, a framework named DISSIMILAR(Megías et al., 2021) was proposed to track and detect false media information by automatically embedding watermarks in images before their release on social media. However, there has been minimal exploration in detecting false content generated by Large Language Models. We propose two potential approaches in this field. The first involves a method similar to the DISSIMILAR framework, where watermarks are added to text content before its publication on social media. This approach would use format-based methods (3.1) to embed watermarks without altering the text’s content. The second approach necessitates collaboration with LLM providers, allowing them to embed watermarks and share watermark detection methods with certain platforms, thereby facilitating the marking of LLM-generated content. We recommend that future work should leverage text watermarking technology to aid in the detection of false information.
@@ 8. Conclusion
This survey has comprehensively explored the landscape of text watermarking in the era of large language models (LLMs). Our investigation covered various aspects, including the implementation text watermarking methods, different perspective of evaluation methods of text watermarking, and applications in areas such as copyright protection, academic integrity, and fake news detection.
One of the critical findings of this survey is the evolution of text watermarking techniques in tandem with the advancements in LLMs. The robustness of watermarking methods against various levels of attacks—character-level, word-level, and document-level—highlights the complexity and sophistication needed in watermark designs to combat increasingly advanced removal strategies. The discussion on unforgeability, especially in the context of private and public detection scenarios, sheds light on the necessity of developing watermarks that are not only difficult to remove but also challenging to replicate or forge.
@@
Future research should focus on developing more advanced watermarking algorithms that can withstand new forms of attacks, especially in scenarios where attackers have access to more sophisticated tools and knowledge. Exploring the use of watermarking in new domains, such as verifying the authenticity of AI-generated content in social media and journalism, can provide new avenues for maintaining the integrity and trustworthiness of digital content.

### v4

@@ 5.4.2. Unforgeability in Publicly Detectable Scenario
Assessing the unforgeability of watermark algorithms under publicly detectable scenarios is considerably more complex since the watermark detector is publicly accessible, providing attackers with a wider range of cracking methods. In such scenarios, attackers could still employ previously mentioned statistical attack methods, such as spoofing attacks (Sadasivan et al., 2023), to analyze and extract the watermark’s generation rules. Additionally, given the close relationship between watermark generation and detection algorithms, attackers may gain insights into the design of the watermark generator by directly analyzing the implementation of the watermark detector.
Currently, the watermark detector in most watermarking algorithms involves details in watermark generator. For example, KGW requires the hash key in watermark generator to determine whether each token is on the green list during the watermark detection process, which exposes the watermark generation method. Consequently, these watermarking algorithms lack unforgeability in publicly detectable scenarios, where the watermark detector is fully accessible to the public. A text watermarking algorithm with unforgeability must ensure that the watermark detector does not reveal information about the watermark generator. For instance, some current methods employ neural networks for watermark detection (Liu et al., 2023b; Abdelnabi and Fritz, 2021), achieving effective detection while preventing further information disclosure due to the black-box nature of neural networks.
@@ 6. Application for Text Watermarking
In preceding sections, we outlined the implementation methods of text watermarking technologies in the era of LLMs and detailed how to thoroughly evaluate these methods. This section delves into their real-world applications, focusing on three areas: copyright protection (Zhao et al., 2023b, 2022; He et al., 2022a, b; Peng et al., 2023; Li et al., 2023; Tang et al., [n. d.]; Chu et al., 2023; Tang et al., 2023; Sun et al., 2023; Wang et al., 2023a; Taleby Ahvanooey et al., 2018; Mir, 2014; Iqbal et al., 2019), academic integrity (Vasilatos et al., 2023; Zhao et al., 2022; Perkins, 2023), and fake news detection (Megías et al., 2021, 2022). First, we analyze the use of text watermarking for copyright protection of large models and datasets, preventing infringement and misuse. Second, we explore its importance in upholding academic integrity, including plagiarism detection and originality verification in academic outputs. Finally, we assess its role in detecting and mitigating fake news, particularly within social media and news platforms. Figure 7 provides an illustrative overview of these applications.
@@ 6.1.1. Text/Dataset Copyright
Text copyright refers to the legal protection of original textual content, such as writings and online posts, ensuring unique rights for the creators. Copyrighted texts should provide sufficient information to identify their creators and sources. This protection extends beyond traditional publications like books and journals to digital-era web articles, blog posts, and other online contents. Present research in textual copyright predominantly centers on format-based watermarking algorithms (see section 3.1), primarily because these algorithms do not necessitate content alteration, which is a critical consideration for some creators.
For example, Taleby Ahvanooey et al. (2018) employs layout attributes like word and line spacing, alongside formatting elements such as text color and font, for watermark insertion. Mir (2014) introduced an invisible digital watermark for web content, using encrypted semantic and syntactic rules converted into spaces with binary control characters, and embedded via HTML structure. Furthermore, specific text formats may require tailored watermarking approaches, as illustrated by Iqbal et al. (2019), who embedded watermarks in MS-Word documents using unique features like variables and bookmarks.
Despite the current reliance on format-based watermarking, the evolution of large language models points to integrating watermark algorithms with these models as a promising avenue for future text copyright protection research.
With the rise and widespread application of deep learning technology, the dataset copyright has become particularly important, which means protecting datasets from unauthorized use has emerged as a crucial issue. Text watermarking technology is central in this domain, creating watermarked datasets by embedding watermarks into data. Models trained on these datasets exhibit identifiable features that discourage unauthorized exploitation.
@@ 6.1.2. LLM Copyright
In the field of copyright protection for LLMs, the key goal is to defend against extraction attacks, where significant data is extracted from LLMs to train new models. A prevalent strategy for copyright protection involves embedding watermarks in LLM outputs. Consequently, attackers inadvertently use watermarked datasets for training, leading to watermarked characteristics in the new model’s outputs. This approach parallels dataset copyright, with the distinction that the watermarked dataset originates from an LLM. Current efforts have developed watermark algorithms for various LLM types, including embedding (Peng et al., 2023), generative (He et al., 2022a, b; Zhao et al., 2023b), and classification (Zhao et al., 2022) LLMs. The input of an embedding LLM is text, and its output is the corresponding embedding of that text. The generative LLM is currently the most commonly used LLM, with both its input and output being text. In the case of a classification LLM, the input is text, and the output is a specific category.
@@ 6.3. Fake News Detection
Large language models (LLMs) pose significant challenges in fake news detection due to their capability to generate and rapidly disseminate false information (Chen and Shu, 2023b, a). The first concern involves their proficiency in creating convincing, yet potentially erroneous or misleading content. This feature makes LLMs effective tools for fabricating fake news, thereby deceiving the public and distorting facts. The second issue is the swift spread of such false information across digital platforms, exacerbating the proliferation of incorrect viewpoints and eroding public trust in reliable sources (DiResta, 2020; Zhou et al., 2023). Thus, identifying LLM-generated news is critical.
Current research in combating fake AI-generated images and videos, such as the DISSIMILAR framework (Megías et al., 2021), employs watermark technology for detection. However, less attention has been paid to false content produced by LLMs. We propose two potential approaches in this field. The first involves a method similar to the DISSIMILAR framework, where watermarks are added to text content before its publication on social media. This approach would use format-based methods (3.1) to embed watermarks without altering the text’s content. The second approach necessitates collaboration with LLM providers, allowing them to embed watermarks and share watermark detection methods with certain platforms, thereby facilitating the marking of LLM-generated content. We recommend that future work should leverage text watermarking technology to aid in the detection of false information.
@@ 8. Conclusion
This survey thoroughly delves into the landscape of text watermarking in the era of large language models (LLMs), encompassing its implementation, evaluation methods, applications in fields like copyright protection, academic integrity, and fake news detection, as well as the challenges and future directions of the domain.
@@
Future research should focus on creating advanced watermarking algorithms capable of withstanding novel attack types, especially where attackers have access to sophisticated tools and knowledge. Exploring watermarking in new applications like authenticity verification of AI-generated content in social media and journalism is crucial for maintaining the integrity and trustworthiness of digital content.